# Robust Neural Information Retrieval: An Adversarial and Out-of-distribution Perspective

### v1

@@ 3.4.4. Empirical defense
Empirical defenses attempt to make models empirically robust to known adversarial attacks; this has been extensively explored in image (Madry et al., 2018) and text classification (Wang et al., 2021). The aim of an empirical defense is to find adversarial examples during training and use them to augment the training set.
@@
In IR, Wu et al. (2022a) have been the first to apply data augmentation to NRM defense. They find that data augmentation can reduce the attack success rate to some extent. However, it performs worse than customized defenses for IR. Hence, simply augmenting the training documents (as in NLP) is not a robust defense against attacks in IR.
@@
Adversarial training is one of the most effective defenses against specific seen attacks. By integrating pre-constructed adversarial samples into training data, the adversarial training has demonstrated a strong defense in both CV and NLP (Madry et al., 2018; Zhu et al., 2019).
@@
To tackle this issue, Liu et al. (2024b) theoretically define the perturbation invariance of a ranking model and design a perturbation-invariant adversarial training (PIAT) method for ranking models to achieve a better effectiveness-robustness trade-off. Experimental results on several ranking models demonstrate the superiority of PITA compared to earlier adversarial defenses.
@@ 3.4.5. Certified robustness
4. Out-of-distribution Robustness
In addition to adversarial vulnerability, deep neural networks lack generalizability to OOD data. When faced with data that differs from the distribution of the training data, neural networks struggle to maintain performance. In the field of IR, this problem has begun to be exposed and attract attention as neural IR models are now widely being used (Wu et al., 2022b; Thakur et al., 2021). Existing work focuses on the OOD generalizability of DRMs, since OOD data has a direct impact on the retrieval stage. We refer back to Section 2.2 for a definition of OOD robustness in IR. In this section, we present specific work of OOD generalizability on unseen documents and OOD generalizability on unseen queries, respectively.
@@ 4.1. Overview
IR systems are widely used in fields such as search engines (Ilan, 1998), digital libraries (Chowdhury and Chowdhury, 2003), medical search (Luo et al., 2008), legal search (Maxwell and Schafer, 2008). At the same time, the scenarios that IR systems need to cope with are becoming more complex. For a long time, research work on neural IR models has often been carried out in a narrow independent and identically distributed (IID) setting. In this setting, the model faces homogeneous data during training and testing.
@@
With deep neural network models being applied to IR, neural IR models have demonstrated excellent results on many tasks (Guo et al., 2020, 2022). However, in the real world, IR systems need to face more than just a single task or scenario (Thakur et al., 2021). In order to deploy an IR system, it is often necessary to construct a training dataset for the neural IR model in it. This process is usually time-consuming and expensive and hence many IR systems are expected to be able to cope with a wide variety of data not seen during training (Thakur et al., 2021; Liu et al., 2023b; Penha et al., 2022). Therefore, OOD generalizability is a key requirement for contemporary IR systems, given the dynamic nature of user needs and evolving data landscapes. In IR, OOD generalizability is focused on unseen documents and unseen queries as shown in Figure 10.
Why should IR models have the generalizability to unseen documents?
@@
Why should IR models have the generalizability to unseen queries?
In practical scenarios, the set of possible queries that users may input is vast and constantly evolving (Penha et al., 2022; Zhuang and Zuccon, 2021). Therefore, IR models must possess the ability to handle queries that were not encountered during training. Without generalizability to unseen queries, IR models risk providing inadequate or irrelevant results, ultimately diminishing user satisfaction and trust in the system (Penha et al., 2022; Campos et al., 2023). The nature of information needs is dynamic and diverse (Guo et al., 2011). Users often express their information needs in varied ways, using different vocabulary, language styles, or even typos (Zhuang et al., 2023; Chen et al., 2022a). A robust neural IR model should be able to understand and accommodate these variations, effectively retrieving relevant information regardless of how the query is formulated. Furthermore, in reality, there is a wide variety of query types that are often not fully or adequately accessible during IR model training (Zhuang and Zuccon, 2022, 2021). But, a good IR system should perform consistently in response to a wide range of query types.
@@ 4.2. Benchmark Datasets
In this section, we present the commonly used datasets for studying OOD robustness in IR. All datasets can be found in the BestIR benchmark.
@@ 4.2.1. Datasets for unseen documents
The datasets for unseen documents mainly involve adaptation to new corpus datasets and updates to corpus datasets. Existing datasets on unseen documents in IR, as shown in Table 2.
Table 2. Benchmark datasets for unseen documents. #Document denotes the number of documents in corpus; #Qtrain denotes the number of queries for training; #Qdev denotes the number of queries for development; #Qeval denotes the number of queries for evaluation; #Corpus denotes the number of corpora.
Type Dataset #Corpus Publication
Adaptation to new corpus BEIR (Thakur et al., 2021) 18 (Oh et al., 2023; Chen et al., 2023a; Anaya-Isaza and Mera-Jiménez, 2022; Laitz et al., 2024; Anand et al., 2023a; Bonifacio et al., 2022; Li and Gaussier, 2024; Ge et al., 2023; Ram et al., 2022; Dai et al., 2022; Liu et al., 2022b; Reddy et al., 2021; Yu et al., 2022; Izacard et al., 2021; Neelakantan et al., 2022; Liang et al., 2020; Ma et al., 2021b; Chandradevan et al., 2024; Thakur et al., 2021; Ma et al., 2021b; Sachan et al., 2023; Cai et al., 2022; Wang et al., 2022b; Zhan et al., 2022; Xian et al., 2023; Xu et al., 2023; Kang et al., 2024; Xin et al., 2022; Yu et al., 2022; Ni et al., 2022; Lu et al., 2022; Huebscher et al., 2022; Formal et al., 2022; Kasela et al., 2024; Chen et al., 2022c; Lee et al., 2023)
Type Dataset #Document #Qtrain #Qdev #Qeval Publication
CDI-MS (Chen et al., 2023d) 3.2M 370K 5,193 5,793 (Chen et al., 2023d)
Updates to a corpus CDI-NQ (Chen et al., 2023d) 8.8M 500K 6,980 6,837 (Chen et al., 2023d)
LL-LoTTE (Cai et al., 2023) 21M 60K 8.8k 3.6k (Cai et al., 2023)
LL-MultiCPR (Cai et al., 2023) 21M 60K 8.8k 3.6k (Cai et al., 2023)
Adaptation to new corpus datasets
In order to model the migration of models between corpora of different domains, adaptation to new corpus datasets typically aggregates multiple existing domain IR datasets. Of these, BEIR (Thakur et al., 2021) is the most typical, it includes nine different retrieval tasks, such as fact-checking, news retrieval, question answering, entity retrieval, and so on. These 18 datasets are across diverse tasks, diverse domains, task difficulties, and diverse annotation strategies.
Updates to a corpus datasets
In order to follow the updates to a corpus over time, the datasets are mainly constructed by slicing or expanding the existing dataset. For example, to mimic the arrival of new documents in MS MARCO (Nguyen et al., 2016), CDI-MS (Chen et al., 2023d) first randomly sampled 60% documents from the whole corpus as the base documents. Then, it randomly samples 10% documents from the remaining corpus as the new document set, which is repeated 4 times. The web is continuously growing with potential distribution shifts. The updates to a corpus focus on the performance of IR model when updates to a corpus undergoes changes over time.
@@ 4.2.2. Datasets for unseen queries
Table 3. Benchmark datasets for unseen queries. #Qeval denotes the number of queries for evaluation.
Type Dataset #Qeval Publication
@@
Query variation datasets contain sets of queries that aim for the same information need but are expressed in various ways, reflecting the natural diversity in how different users might phrase their search queries. Such datasets can include paraphrased queries, queries with typos, order-swapped queries, and queries without stop words. For example, Penha et al. (2022) construct query variation datasets by turning queries from TREC DL19 (Craswell et al., 2020b) and ANTIQUE (Hashemi et al., 2020) into different variants using 4 categories in 10 ways. The importance of query variation datasets lies in their ability to simulate real-world search scenarios, where users often have unique ways of expressing their information needs.
@@
4.3. OOD generalizability on unseen documents
As argued above, IR systems need to adapt to different environments and variations in the corpus. However, retraining the neural IR models in each new environment is costly. Previous work has only analyzed the generalizability of existing IR models across different domains (Wu et al., 2022b; Thakur et al., 2021; Ren et al., 2023). In this work, we summarize work on adaptation to a new corpus and updates to a corpus. Figure 11 illustrates how we organize the discussion of different methodologies.
OOD generalizability on unseen documents Updates to a corpus Continual learning for dense retrieval (Cai et al., 2023) Continual learning for generative retrieval (Yoon et al., 2023; Kishore et al., 2023; Chen et al., 2023d; Guo et al., 2024; Mehta et al., 2023a) Adaptation to new corpus Scaling up the model capacity (Ni et al., 2022; Lu et al., 2022) Architectural modifications (Huebscher et al., 2022; Formal et al., 2022; Kasela et al., 2024; Chen et al., 2022c; Lee et al., 2023) Domain modeling (Zhan et al., 2022; Xian et al., 2023; Xu et al., 2023; Kang et al., 2024; Xin et al., 2022; Yu et al., 2022) Data augmentation (Oh et al., 2023; Chen et al., 2023a; Anaya-Isaza and Mera-Jiménez, 2022; Laitz et al., 2024; Anand et al., 2023a; Bonifacio et al., 2022; Li and Gaussier, 2024; Ge et al., 2023; Ram et al., 2022; Dai et al., 2022; Liu et al., 2022b; Reddy et al., 2021; Yu et al., 2022; Izacard et al., 2021; Neelakantan et al., 2022; Liang et al., 2020; Ma et al., 2021b; Chandradevan et al., 2024; Thakur et al., 2021; Ma et al., 2021b; Sachan et al., 2023; Cai et al., 2022; Wang et al., 2022b)
@@ 4.3.1. Definition
Adaptation to new corpora refers to the trained IR models that hard to adapt to a corpus of new domains in the absence of supervised data (Thakur et al., 2021). The goal of improving generalizability on the adaptation to new corpus for a neural IR model $f$ under top-$K$ ranked results can usually be formalized as:
@@
where $o$ is the original corpus domain and $n$ is the new corpus domain. Among the main solutions for adaptation to a new corpus are data augmentation (Wang et al., 2022b; Izacard et al., 2021; Bonifacio et al., 2022) , domain modeling (Yu et al., 2022), architectural modifications (Kasela et al., 2024), and scaling up the model capacity (Ni et al., 2022).
@@
Updates to a corpus refer to the problem for the trained IR model cannot maintain the ranking performance under continuously arrived new documents (Chen et al., 2023d). The goal of improving generalizability on the updates to a corpus for a neural IR model $f$ under top-$K$ ranked results can be formalized as:
(16) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}^{\sum^{t}_{0}}};\mathcal{D}_{\mathrm{test}}^{t+1}\right)\text{ such that }\mathcal{C}_{\sim\mathcal{G}}^{\sum^{t}_{0}}\in\mathcal{D}_{\mathrm{train}}^{\sum^{t}_{0}},\;\mathcal{C}_{\sim\tilde{\mathcal{G}}}^{t+1}\in\mathcal{D}_{\mathrm{test}}^{t+1}.$
@@
In the following, we first introduce the evaluation metrics widely used for OOD generalizability on unseen documents. Then, we detail the existing solutions for adaptation to a new corpus and updates to a corpus, respectively.
@@ 4.3.2. Evaluation
OOD generalizability of IR models on unseen documents is mainly measured by the ranking performance under the new corpus. For both adaptation to a new corpus and updates to a corpus, ranking performance is the common evaluation. For updates to a corpus, specifically considering the evaluation of the degree of forgetting of the old corpus.
@@
For adaptation to a new corpus and updates to a corpus, the ranking performance of IR models under unseen documents is evaluated by common metrics.
  • NDCG (Järvelin and Kekäläinen, 2002) evaluates the quality of ranked results by measuring the gain of a document based on its position in the ranked list.
  • MRR (Craswell, 2009) evaluates the performance of a ranking result by calculating the average of the reciprocal ranks of the first relevant document answer.
  • HIT (Chen et al., 2023d) evaluates the proportion of times a relevant document is found within a set of top-N ranking results.
  • AP (Mehta et al., 2023b) evaluates the average performance of the ranking performance metrics, overall new domains in adaptation to new corpus, and sessions in updates to a corpus. The ranking performance metric could be any of the above.
Metrics for forgetting degree of the old corpus
updates to a corpus are an ongoing process with many sessions and requires the model to memorize new data without forgetting the old. Therefore, some metrics have been proposed to evaluate the model performance from a time-series perspective.
  • Training Time (Cai et al., 2023) evaluates the total time it takes for the IR model to learn new data while recalling old data.
@@
    where $p$ is the ranking performance under any common metrics.

### v2

@@ 3.4.4. Empirical defense
Empirical defenses attempt to make models empirically robust to known adversarial attacks; this has been extensively explored in image retrieval (Madry et al., 2018) and text classification (Wang et al., 2021). The aim of an empirical defense is to find adversarial examples during training and use them to augment the training set.
@@
In IR, Wu et al. (2022a) have been the first to apply data augmentation to the defense of NRMs. They find that data augmentation can reduce the attack success rate to some extent. However, it performs worse than customized defenses for IR. Hence, simply augmenting the training documents (as in NLP) is not a robust defense against attacks in IR.
@@
Adversarial training is one of the most effective defenses against specific seen attacks. By integrating pre-constructed adversarial samples into training data, adversarial training has demonstrated a strong defense in both CV and NLP (Madry et al., 2018; Zhu et al., 2019).
@@
To tackle this issue, Liu et al. (2024b) theoretically define the perturbation invariance of a ranking model and design a perturbation-invariant adversarial training (PIAT) method for ranking models to achieve a better effectiveness-robustness trade-off. Experimental results on several ranking models demonstrate the superiority of PIAT compared to earlier adversarial defenses.
@@ 3.4.5. Certified robustness
4. OOD Robustness
In addition to adversarial vulnerability, deep neural networks lack generalizability to OOD data. When faced with data that differs from the distribution of the training data, neural networks struggle to maintain performance. In IR, this problem has begun to be exposed and attract attention as neural IR models are now widely being used (Wu et al., 2022b; Thakur et al., 2021). Prior work focuses on the OOD generalizability of DRMs, since OOD data has a direct impact on the retrieval stage. We refer back to Section 2.2 for a definition of OOD robustness in IR. In this section, we present specific work on OOD generalizability on unseen documents and OOD generalizability on unseen queries, respectively.
Adapt to OOD corpusAdapt to OOD query setNeural IR modelTraining corpusTraining queries
@@ 4.1. Overview
For a long time, research work on neural IR models has been carried out in a narrow independent and identically distributed (IID) setting. In this setting, the model faces homogeneous data during training and testing. But IR systems are widely used in fields such as search engines (Ilan, 1998), digital libraries (Chowdhury and Chowdhury, 2003), medical search (Luo et al., 2008), legal search (Maxwell and Schafer, 2008), and at the same time, the scenarios that IR systems need to cope with are becoming more complex.
@@
With deep neural network models being applied to IR, neural IR models have demonstrated excellent results on many tasks (Guo et al., 2020, 2022). But IR systems need to face more than just a single task or scenario (Thakur et al., 2021). In order to deploy an IR system, it is often necessary to construct a training dataset for the neural IR model in it. This process is usually time-consuming and expensive and hence many IR systems are expected to be able to cope with a wide variety of data not seen during training (Thakur et al., 2021; Liu et al., 2023b; Penha et al., 2022). Therefore, OOD generalizability is a key requirement for contemporary IR systems, given the dynamic nature of user needs and evolving data landscapes. In IR, OOD generalizability is focused on unseen documents and unseen queries as illustrated in Figure 9.
Why should IR models be able to generalize to unseen documents?
@@
Why should IR models be able to generalize to unseen queries?
The set of possible queries that users may input is vast and constantly evolving (Graus et al., 2018; Penha et al., 2022; Zhuang and Zuccon, 2021). E.g., 15% of daily Google searches are brand new. [^10] https://blog.google/products/search/our-latest-quality-improvements-search/ The nature of information needs is dynamic and diverse (Guo et al., 2011; Lefortier et al., 2014). Users often express their information needs in varied ways, using different vocabulary, language styles, or even typos (Zhuang et al., 2023; Chen et al., 2022a). This challenge becomes particularly pronounced in the context of ever-changing user interests and the introduction of new vocabularies. Therefore, IR models must possess the ability to handle queries that were not encountered during training. Without generalizability to unseen queries, IR models risk providing inadequate or irrelevant results, ultimately diminishing user satisfaction and trust in the system (Penha et al., 2022; Campos et al., 2023). A robust neural IR model should be able to understand and accommodate these variations, effectively retrieving relevant information regardless of how the query is formulated.
Furthermore, in reality, there is a wide variety of query types that are often not fully or adequately accessible during IR model training (Zhuang and Zuccon, 2022, 2021). But, a robust IR system should perform consistently in response to a wide range of query types.
@@ 4.2. Benchmark Datasets
In this section, we present the commonly used datasets for studying OOD robustness in IR. All datasets can be found in the BestIR benchmark; details about BestIR can be found in Appendix B.
@@ 4.2.1. Datasets for unseen documents
The datasets for unseen documents mainly involve adaptation to new corpus datasets and updates to corpus datasets. Datasets on unseen documents in IR are listed in Table 2.
Table 2. Benchmark datasets for unseen documents. #Documents denotes the number of documents in corpus; #Qtrain denotes the number of queries available for training; #Qdev denotes the number of queries available for development; #Qeval denotes the number of queries available for evaluation; #Corpora denotes the number of corpora.
Type Dataset #Corpora Publications
Adaptation to a new corpus BEIR (Thakur et al., 2021) 18 (Oh et al., 2023; Chen et al., 2023a; Anaya-Isaza and Mera-Jiménez, 2022; Laitz et al., 2024; Anand et al., 2023a; Bonifacio et al., 2022; Li and Gaussier, 2024; Ge et al., 2023; Ram et al., 2022; Dai et al., 2022; Liu et al., 2022b; Reddy et al., 2021; Yu et al., 2022; Izacard et al., 2021; Neelakantan et al., 2022; Liang et al., 2020; Ma et al., 2021b; Chandradevan et al., 2024; Thakur et al., 2021; Ma et al., 2021b; Sachan et al., 2023; Cai et al., 2022; Wang et al., 2022b; Zhan et al., 2022; Xian et al., 2023; Xu et al., 2023; Kang et al., 2024; Xin et al., 2022; Yu et al., 2022; Ni et al., 2022; Lu et al., 2022; Huebscher et al., 2022; Formal et al., 2022; Kasela et al., 2024; Chen et al., 2022c; Lee et al., 2023)
Type Dataset #Documents #Qtrain #Qdev #Qeval Publications
Updates to a corpus CDI-MS (Chen et al., 2023d) 3.2M 370K 5,193 5,793 (Chen et al., 2023d)
CDI-NQ (Chen et al., 2023d) 8.8M 500K 6,980 6,837 (Chen et al., 2023d)
LL-LoTTE (Cai et al., 2023) 5.5M 16K 8.5K 8.6K (Cai et al., 2023)
LL-MultiCPR (Cai et al., 2023) 3.0M 136K 15K 15K (Cai et al., 2023)
Adaptation to a new corpus
In order to model the migration of models between corpora of different domains, datasets that cater for adaptation to a new corpus typically aggregate multiple domain-specific IR datasets. Of these, BEIR (Thakur et al., 2021) is the best-known example; it includes nine retrieval tasks, such as fact-checking, news retrieval, question answering, entity retrieval. It also has 18 datasets, across diverse tasks, diverse domains, task difficulties, and diverse annotation strategies.
Updates to a corpus
Updates to a corpus focus on the performance of an IR model when updates to a corpus occur. In order to follow updates to a corpus over time, the available datasets are mainly constructed by slicing or expanding the existing dataset. For example, to mimic the arrival of new documents in MS MARCO (Nguyen et al., 2016), CDI-MS (Chen et al., 2023d) first randomly sampled 60% documents from the whole corpus as the base documents. Then, it randomly samples 10% documents from the remaining corpus as the new document set, which is repeated 4 times.
@@ 4.2.2. Datasets for unseen queries
Table 3. Benchmark datasets for unseen queries. #Qeval denotes the number of queries available for evaluation.
Type Dataset #Qeval Publications
@@
The importance of query variation datasets lies in their ability to simulate real-world search scenarios, where users often have unique ways of expressing their information needs. Query variation datasets contain sets of queries that target the same information need but are expressed in alternative ways, reflecting the natural diversity in how different users might phrase their search queries. Such datasets can include paraphrased queries, queries with typos, order-swapped queries, and queries without stop words. For example, Penha et al. (2022) construct query variation datasets by turning queries from TREC DL19 (Craswell et al., 2020b) and ANTIQUE (Hashemi et al., 2020) into different variants using 4 categories in 10 ways.
@@
4.3. OOD Generalizability to Unseen Documents
As argued above, IR systems need to adapt to different environments and variations in the corpus. However, retraining the neural IR models in each new environment is costly. Previous work has only analyzed the generalizability of IR models across different domains (Wu et al., 2022b; Thakur et al., 2021; Ren et al., 2023). In this work, we summarize work on adaptation to a new corpus and updates to a corpus. Figure 10 illustrates how we organize the discussion of different methodologies.
OOD generalizability on unseen documents Updates to a corpus Continual learning for dense retrieval (Cai et al., 2023) Continual learning for generative retrieval (Yoon et al., 2023; Kishore et al., 2023; Chen et al., 2023d; Guo et al., 2024; Mehta et al., 2023a) Adaptation to a new corpus Scaling up the model capacity (Ni et al., 2022; Lu et al., 2022) Architectural modifications (Huebscher et al., 2022; Formal et al., 2022; Kasela et al., 2024; Chen et al., 2022c; Lee et al., 2023) Domain modeling (Zhan et al., 2022; Xian et al., 2023; Xu et al., 2023; Kang et al., 2024; Xin et al., 2022; Yu et al., 2022) Data augmentation (Oh et al., 2023; Chen et al., 2023a; Anaya-Isaza and Mera-Jiménez, 2022; Laitz et al., 2024; Anand et al., 2023a; Bonifacio et al., 2022; Li and Gaussier, 2024; Ge et al., 2023; Ram et al., 2022; Dai et al., 2022; Liu et al., 2022b; Reddy et al., 2021; Yu et al., 2022; Izacard et al., 2021; Neelakantan et al., 2022; Liang et al., 2020; Ma et al., 2021b; Chandradevan et al., 2024; Thakur et al., 2021; Ma et al., 2021b; Sachan et al., 2023; Cai et al., 2022; Wang et al., 2022b)
@@ 4.3.1. Definition
Adaptation to a new corpus refers to the trained IR models that may be hard to adapt to a corpus of new domains in the absence of supervised data (Thakur et al., 2021). The goal of improving generalizability on the adaptation to a new corpus for a neural IR model $f$ under top-$K$ ranked results can usually be formalized as:
@@
where $o$ is the domain of the original corpus and $n$ is the new corpus domain. Among the main solutions for adaptation to a new corpus are data augmentation (Wang et al., 2022b; Izacard et al., 2021; Bonifacio et al., 2022), domain modeling (Yu et al., 2022), architectural modifications (Kasela et al., 2024), and scaling up the model capacity (Ni et al., 2022).
@@
Updates to a corpus refer to the problem for the trained IR model to maintain its ranking performance under continuously arriving new documents (Chen et al., 2023d). The goal of improving generalizability on the updates to a corpus for a neural IR model $f$ under top-$K$ ranked results can be formalized as:
(16) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}^{\sum^{t}_{0}}};\mathcal{D}_{\mathrm{test}}^{t+1}\right)\text{ such that }\mathcal{C}_{\sim\mathcal{G}}^{\sum^{t}_{0}}\in\mathcal{D}_{\mathrm{train}}^{\sum^{t}_{0}},\;\mathcal{C}_{\sim\tilde{\mathcal{G}}}^{t+1}\in\mathcal{D}_{\mathrm{test}}^{t+1},$
@@
Below, we first introduce the evaluation metrics widely used for OOD generalizability on unseen documents. Then, we detail solutions for adaptation to a new corpus and updates to a corpus, respectively.
@@ 4.3.2. Evaluation
OOD generalizability of IR models on unseen documents is mainly measured by the ranking performance under the new corpus. For both adaptation to a new corpus and updates to a corpus, ranking performance is the common evaluation. For updates to a corpus, previous work also evaluates the degree to which the old corpus is forgotten.
@@
For adaptation to a new corpus and updates to a corpus, the ranking performance of IR models under unseen documents is evaluated by common metrics:
  • NDCG (Järvelin and Kekäläinen, 2002) evaluates the quality of ranked results by measuring the gain of a document based on its position in the ranked list;
  • MRR (Craswell, 2009) evaluates the performance of a ranking result by calculating the average of the reciprocal ranks of the first relevant document answer;
  • HIT (Chen et al., 2023d) evaluates the proportion of times a relevant document is found within a set of top-$N$ ranked results; and
  • AP (Mehta et al., 2023b) evaluates the average performance of the ranking performance metrics, overall new domains in adaptation to new corpus, and sessions in updates to a corpus; the ranking performance metric could be any of the above.
Metrics for the degree of forgetting the old corpus
Updates to a corpus are an ongoing process with many sessions; they require that the model memorizes new data without forgetting the old. Therefore, some metrics have been proposed to evaluate the model performance from a time-series perspective.
  • Training Time (Cai et al., 2023) evaluates the total time it takes for the IR model to learn new data while recalling old data;
@@
    where $p$ is the ranking performance under any common metrics; and